Compute the Internet checksum for the data block E3 4F 2A 8A 44 26 86 4F. Then check for verification.
Answer #1

Given data block: E34F 2A8A 4426 864F

Sum of given data: E34F+2A8A+4426+864F = D84F

Take complement of D84F to obtain Checksum = (FFFF-D84F) = 27B0

Message with the checksum will be = E34F 2A8A 4426 864F 27B0

To verify add all: E34F+2A8A+4426+864F+27B0 = FFFF (no error/message is correct)